A hammer falls from a shelf and strikes the ground at 12 ms-1
. Using kinetic
and potential energies in the calculation, determine what height did it fell from.
1
Expert's answer
2020-11-23T10:27:42-0500

We can find the height from which the hammer was fell from the Law of Conservation of Energy:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""h=\dfrac{v^2}{2g}=\dfrac{(12\ \dfrac{m}{s})^2}{2\cdot 9.8\ \dfrac{m}{s^2}}=7.35\ m."

Answer:

"h=7.35\ m."